# FAB-PPI: Frequentist, Assisted by Bayes, Prediction-Powered Inference

### v1

@@ Acknowledgements
$\Pr(\Delta_{\theta}\in\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))=\Pr\left(-(z_{1-\delta}-\epsilon)<\frac{\widehat{\Delta}_{\theta}-\Delta_{\theta}}{\widehat{\sigma}_{\theta}}<\frac{2\Delta_{\theta}}{\widehat{\sigma}_{\theta}}+z_{1-\delta}-\epsilon\right).$ (S34)
@@
${\lim}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq{\lim}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq 1-\delta.$ (S35)
Lemma S3.2.
For $\delta\in(0,1)$, let $w_{\delta}(\cdot;\widehat{\sigma}_{\theta})$ be the FAB spending function associated to the FAB confidence region
$\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\bullet)=\text{FAB-CR}(\bullet;\pi_{0}(\cdot~{};\widehat{\sigma}_{\theta}),\widehat{\sigma}_{\theta},\delta).$
Assume $\widehat{\sigma}_{\theta}\to 0$ almost surely as $\min(n,N)\to\infty$. If $\pi_{0}$ is the horseshoe prior (23), then, for any $z\in\mathbb{R}$,
$w_{\delta}(z;\widehat{\sigma}_{\theta})\to\frac{1}{2}\text{ a.s. as }\min(n,N)\to\infty.$
Proof.
As described in Section S1.2, the spending function $w_{\delta}$, which is continuous, satisfies, for any $z\in\mathbb{R}$ and $\sigma>0$,
$w_{\delta}\left(z;\sigma\right)=w_{\delta}\left(\frac{z}{\sigma};1\right).$
By Proposition S1.3, we have
Since $\widehat{\sigma}_{\theta}\to 0$ almost surely and $w_{\delta}$ is continuous, for any $z\in\mathbb{R}$, we conclude that
$w_{\delta}\left(z;\widehat{\sigma}_{\theta}\right)\to\frac{1}{2}\text{ a.s. as}\min(n,N)\to\infty.$
@@
as $\min(n,N)\to\infty$, where $\frac{\widehat{\sigma}_{\theta}^{2}}{\mathrm{var}(\widehat{\Delta}_{\theta})}\to 1$ almost surely. Specify a symmetric prior $\pi_{0}(\cdot;\widehat{\sigma}_{\theta})$ for $\Delta_{\theta}$ and consider the corresponding $1-\delta$ FAB confidence region
@@
with associated weight function $w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})$. If, for any $x\in\mathbb{R}$,
$w_{\delta}\left(x;\widehat{\sigma}_{\theta}\right)\to\frac{1}{2}\text{ a.s. as }\min(n,N)\to\infty,$ (S36)
then the confidence region $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ reverts to the classical $1-\delta$ $z$-interval for $\Delta_{\theta}$, i.e., almost surely,
$\lim_{\min(n,N)\to\infty}\frac{\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))-\widehat{\Delta}_{\theta}}{\widehat{\sigma}_{\theta}}=\lim_{\min(n,N)\to\infty}\frac{\widehat{\Delta}_{\theta}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))}{\widehat{\sigma}_{\theta}}=z_{1-\delta/2},$
with
@@
As $\Phi^{-1}$ is continuous, Equation S36 implies that, for any $x\in\mathbb{R}$,
$\lim_{\min(n,N)\to\infty}z_{1-\delta w_{\delta}(x;\widehat{\sigma}_{\theta})}=\lim_{\min(n,N)\to\infty}z_{1-\delta(1-w_{\delta}(x;\widehat{\sigma}_{\theta}))}=z_{1-\delta/2}$
almost surely. From Equation S32, it follows that, for any $\epsilon>0$ and for $\min(n,N)$ large enough,
@@
$z_{1-\delta/2}-\epsilon\leq\frac{\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))-\widehat{\Delta}_{\theta}}{\widehat{\sigma}_{\theta}}\leq z_{1-\delta/2}+\epsilon$
$z_{1-\delta/2}-\epsilon\leq\frac{\widehat{\Delta}_{\theta}-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))}{\widehat{\sigma}_{\theta}}\leq z_{1-\delta/2}+\epsilon.$
@@
(S37) then follows directly, which completes the proof. ∎
@@
[…]
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta};\widehat{\sigma}_{\theta}^{f})$ $=\left(\widehat{m}_{\theta}-\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2},\widehat{m}_{\theta}+\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2}\right),$
@@
has correct asymptotic coverage, i.e. it satifies
$\liminf_{\min(n,N)\to\infty}\Pr(\theta^{*}\in\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}})=\liminf_{\min(n,N)\to\infty}\Pr(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}}^{f}))\geq 1-\alpha.$
@@
By Lemma S3.1 (Gaussian prior) and Lemmas S3.2 and S3.3 (horseshoe prior), $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})$ is an asymptotically valid $1-\delta$ confidence region for $\Delta_{\theta^{\star}}$, that is
$\liminf_{n,N\to\infty}\text{Pr}(\Delta_{\theta^{\star}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}))$ $\geq 1-\delta.$
@@
$\liminf_{n,N\to\infty}\text{Pr}(m_{\theta^{*}}\in\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f}))$ $\geq 1-(\alpha-\delta).$
@@
$E=\{\Delta_{\theta^{*}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})\}\cap\{m_{\theta^{*}}\in\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f})\}.$
@@
$\liminf_{n,N\to\infty}\mathrm{Pr}(E)$ $\geq 1-\limsup_{n,N\to\infty}\mathrm{Pr}(\{\Delta_{\theta^{*}}\notin\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})\}\cup\{m_{\theta^{*}}\notin\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f})\})$
$\geq 1-\limsup_{n\to\infty}\mathrm{Pr}(\{\Delta_{\theta^{*}}\notin\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})\})-\limsup_{N\to\infty}\mathrm{Pr}(\{m_{\theta^{*}}\notin\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f})\})$
@@
$0=\Delta_{\theta^{*}}+m_{\theta^{*}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f}),$
@@
$\liminf_{n,N\to\infty}\text{Pr}(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{*}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{*}};\widehat{\sigma}_{\theta^{*}}^{f}))\geq 1-\alpha,$
@@
S3.2 Proposition 4.2 - Robustness of FAB-PPI under the horseshoe prior
@@
We first state a corollary of Cortinovis & Caron (2024, Theorem 2.2), which follows from the power-law tails of the marginal likelihood under the horseshoe prior (see Section S1.1). The corollary states that, if $|\widehat{\Delta}_{\theta}|$ is very large, then the standard CLT-based confidence interval, $(\widehat{\Delta}_{\theta}\pm\widehat{\sigma}_{\theta}z_{1-\delta/2})$, is recovered.
@@
$\lim_{|\Delta|\to\infty}\sup(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\Delta;\sigma))-\Delta$ $=\lim_{|\Delta|\to\infty}\Delta-\inf(\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\Delta;\sigma))=\sigma z_{1-\delta/2}.$
@@
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta};\widehat{\sigma}_{\theta}^{f})=\left(\widehat{m}_{\theta}-\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2},\widehat{m}_{\theta}+\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2}\right)$
@@
$\lim_{|\Delta|\to\infty}\sup(\mathcal{S}_{\alpha,\delta}(\Delta,\sigma,m,\sigma^{f}))-(\Delta+m)$ $=\lim_{|\Delta|\to\infty}(\Delta+m)-\inf(\mathcal{S}_{\alpha,\delta}(\Delta,\sigma,m,\sigma^{f}))=\sigma z_{1-\delta/2}+\sigma^{f}z_{1-(\alpha-\delta)/2}.$
Therefore, if $|\widehat{\Delta}_{\theta}|\gg 0$, the confidence region $\mathcal{S}_{\alpha,\delta}(\widehat{\Delta}_{\theta},\widehat{\sigma}_{\theta},\widehat{m}_{\theta},\widehat{\sigma}_{\theta}^{f})$ reverts to the standard interval
$(\widehat{\Delta}_{\theta}+\widehat{m}_{\theta}\pm(\widehat{\sigma}_{\theta}z_{1-\delta/2}+\widehat{\sigma}^{f}_{\theta}z_{1-(\alpha-\delta)/2})).$
It follows that, if $\inf_{\theta^{\prime}\in\mathbb{R}}|\widehat{\Delta}_{\theta^{\prime}}|\gg 0$ the confidence region for $\theta^{\star}$, defined as
@@
S3.3 Proposition 4.3 - Consistency of FAB-PPI mean estimators
@@
[…]
@@
Under the horseshoe prior (23), $\ell_{\text{HS}}^{\prime}(y;1,1)$ is bounded. Therefore (S39) and (S40) hold almost surely by sandwiching.
@@
as $\min(n,N)\to\infty$, where $\widehat{\Sigma}_{\theta}$ is an estimator of $\mathrm{cov}(\widehat{\Delta}_{\theta})$. Again, this holds for both the PPI and PPI++ estimators (Angelopoulos et al., 2023b, a). We consider $d$ independent priors, $\pi_{0}(\Delta_{\theta,k};\widehat{\sigma}_{\theta,k})$ for $k=1,\dots,d$, on the components of $\Delta_{\theta}$, where $\widehat{\sigma}_{\theta,k}^{2}$ is the $k$-th diagonal element of $\widehat{\Sigma}_{\theta}$. The multivariate FAB-PPI estimator $\widehat{\Delta}_{\theta}^{\mathrm{\texttt{FABPP}}}$ is formed by stacking the individual estimators
$\widehat{\Delta}_{\theta,k}^{\mathrm{\texttt{FABPP}}}=\widehat{\Delta}_{\theta,k}+\widehat{\sigma}_{\theta,k}^{2}\ell^{\prime}\left(\widehat{\Delta}_{\theta,k};\widehat{\sigma}_{\theta,k}\right)$
@@
exactly as in the univariate case. Moreover, also multivariate FAB-PPI enjoys exact asymptotic coverage as $\min(n,N)\to\infty$. In particular, Theorem 4.1 can be easily extended to the multivariate case by applying a union bound over the dimensions of $\Delta_{\theta}$.
@@
Here we provide a brief description of each dataset used for the mean estimation experiments in Section 5.2. For additional details, the reader may refer to Angelopoulos et al. (2023b). All of the datasets were downloaded from the examples provided as part of the ppi-py package (Angelopoulos et al., 2023a).

### v2

@@ Acknowledgements
$\Pr(\Delta_{\theta}\in\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))=\Pr\left(-(z_{1-\delta}-\epsilon)\leq\frac{\widehat{\Delta}_{\theta}-\Delta_{\theta}}{\widehat{\sigma}_{\theta}}\leq\frac{2\Delta_{\theta}}{\widehat{\sigma}_{\theta}}+z_{1-\delta}-\epsilon\right).$ (S35)
@@
${\lim\inf}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq{\lim\inf}_{\min(n,N)\to\infty}\Pr(\Delta_{\theta}\in\mathcal{S}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})\mid\Delta_{\theta})\geq 1-\delta.$ (S36)
The proof proceeds similarly for $\Delta_{\theta}<0$. ∎
@@
as $\min(n,N)\to\infty$, where $\widehat{\sigma}_{\theta}^{2}/\mathrm{var}(\widehat{\Delta}_{\theta})\to 1$ almost surely. Let the prior $\pi_{0}(\cdot;\widehat{\sigma}_{\theta})$ on $\Delta_{\theta}$ be the horseshoe prior (23) with scale parameter $\widehat{\sigma}_{\theta}$ and consider the corresponding $1-\delta$ FAB confidence region
@@
where $w_{\delta}(\Delta_{\theta};\widehat{\sigma}_{\theta})$ is the associated weight function. Then, for $\Delta_{\theta}\neq 0$, the confidence region $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})$ reverts to the classical $1-\delta$ $z$-interval for $\Delta_{\theta}$, i.e., almost surely,
$\lim_{\min(n,N)\to\infty}\frac{\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})-\widehat{\Delta}_{\theta}}{\widehat{\sigma}_{\theta}}=[-z_{1-\delta/2},z_{1-\delta/2}],$
where the convergence is with respect to the Hausdorff distance on closed subsets of $\mathbb{R}$. Moreover, for any $\Delta_{\theta}\in\mathbb{R}$,
@@
Consider the case $\Delta_{\theta}\neq 0$. As described in Section S1.2, the spending function $w_{\delta}$ is continuous and satisfies, for any $z\in\mathbb{R}$ and $\sigma>0$,
$w_{\delta}\left(z;\sigma\right)=w_{\delta}\left(\frac{z}{\sigma};1\right).$ (S38)
Moreover, by Section S1.2, $w_{\delta}(\cdot;1)$ takes values in $(0,1)$ and satisfies
Define $A(p)=-\Phi^{-1}(1-\delta(1-p))$ and $B(p)=\Phi^{-1}(1-\delta p)$, where $\Phi(\cdot)$ is the CDF of the standard normal distribution. Then, from Equation S33, we have that
@@
$\frac{\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})-\widehat{\Delta}_{\theta}}{\widehat{\sigma}_{\theta}}$ $=\left\{\psi\in\mathbb{R}\mid A(w_{\delta}(\widehat{\sigma}_{\theta}\psi+\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))\leq\psi\leq B(w_{\delta}(\widehat{\sigma}_{\theta}\psi+\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta}))\right\}$
$=\left\{\psi\in\mathbb{R}\mid A(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))\leq\psi\leq B(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))\right\}$
$=:\mathcal{C}_{n,N},$
where the second equality follows from Equation S38.
Assume that $\Delta_{\theta}>0$, which ensures $\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta}\to\infty$ almost surely as $\min(n,N)\to\infty$. The case $\Delta_{\theta}<0$ follows similarly.
First, we show that there exists an $M$ independent of $n,N$ such that, for all $n,N\geq 1$, if $\psi\in\mathcal{C}_{n,N}$, then $\psi>M$. By the boundedness of $w_{\delta}(\cdot;1)$, there exists $\kappa\in(0,\frac{1}{2})$ such that $w_{\delta}(x;1)\in[\kappa,1-\kappa]$ for all $x\in\mathbb{R}$. Since $A(\cdot)$ is decreasing,
$A(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))\geq A(1-\kappa):=c>$-$\infty$
for all $\psi\in\mathbb{R}$, $n,N\geq 1$. Pick $M<c$. For all $\psi\leq M$ we have, for all $n,N\geq 1$, $\psi\leq M<c\leq A(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))$. Hence, $\psi\notin\mathcal{C}_{n,N}$. As a result, $\mathcal{C}_{n,N}\cap[M,\infty)=\mathcal{C}_{n,N}$.
Second, we show that, almost surely, $w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1)$ converges to $1/2$ uniformly on $M,\infty)$ as $\min(n,N)\to\infty$. Given that $w_{\delta}(\cdot;1)$ is bounded and converges pointwise to $1/2$, we have that
$g(t):=\sup_{y\geq t}\left|w_{\delta}(y;1)-\frac{1}{2}\right|\to 0$
as $t\to\infty$, where $g(\cdot)$ is continuous and nonincreasing. As a result of this, and since $\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta}\rightarrow\infty$ almost surely as $\min(n,N)\to\infty$, we have that
$\sup_{\psi\geq M}\left|w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1)-\frac{1}{2}\right|=g\left(\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta}+M\right)\to 0$ (S39)
almost surely as $\min(n,N)\to\infty$.
Lastly, we combine the previous two steps to show the almost sure Hausdorff convergence of $\mathcal{C}_{n,N}$. The functions $A(\cdot)$ and $B(\cdot)$ are continuous with $A(1/2)=-z_{1-\delta/2}$ and $B(1/2)=z_{1-\delta/2}$. Hence, for every $\epsilon\in(0,z_{1-\delta/2})$, there exists $\eta>0$ such that
$\left|p-\frac{1}{2}\right|\leq\eta\Rightarrow\left|A(p)+z_{1-\delta/2}\right|\leq\epsilon\text{ and }\left|B(p)-z_{1-\delta/2}\right|\leq\epsilon$
By Equation S39, there exists $N_{1}$ such that, for all $\min(n,N)\geq N_{1}$,
$\sup_{\psi\geq M}\left|w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1)-\frac{1}{2}\right|\leq\eta.$
Then, uniformly for $\psi\geq M$,
$A(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))$ $\in[-z_{1-\delta/2}-\epsilon,-z_{1-\delta/2}+\epsilon]$
$B(w_{\delta}(\psi+\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta};1))$ $\in[z_{1-\delta/2}-\epsilon,z_{1-\delta/2}+\epsilon]$
and, for $\min(n,N)\geq N_{1},$
$\left[-z_{1-\delta/2}+\epsilon,z_{1-\delta/2}-\epsilon\right]\subseteq\mathcal{C}_{n,N}\cap[M,\infty)\subseteq[-z_{1-\delta/2}-\epsilon,z_{1-\delta/2}+\epsilon],$
@@
which, combined with the first step, gives the desired result. Moreover, Equation S37 then follows directly, which completes the proof.
On the other hand, in the case $\Delta_{\theta}=0$, asymptotic coverage follows like in the proof of Section S3.1. In particular, from Equation S33, we have that
$\Pr(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta};\widehat{\sigma}_{\theta})|\Delta_{\theta}=0)=\Pr(|\widehat{\Delta}_{\theta}|\leq\widehat{\sigma}_{\theta}z_{1-\delta/2})$
thanks to the fact $w_{\delta}(0;\sigma)=1/2$, and the result follows from the CLT as $\min(n,N)\to\infty$. While this is not necessary to show asymptotic coverage, it is interesting to note that $\mathcal{C}_{n,N}$ does not converge to a deterministic limit almost surely when $\Delta_{\theta}=0$. Instead, by showing continuity of the mapping $\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta}\mapsto\mathcal{C}_{n,N}(\widehat{\Delta}_{\theta}/\widehat{\sigma}_{\theta})$ and again exploiting the CLT, one may prove that $\mathcal{C}_{n,N}$ converges in distribution to the random FAB confidence region $\mathcal{C}(Y)-Y$, where $\mathcal{C}(y)$ is defined in Equation S29, for a unit scale and a horseshoe prior, and $Y\sim\mathcal{N}(0,1)$. ∎
@@
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta};\widehat{\sigma}_{\theta}^{f})$ $=\left[\widehat{m}_{\theta}\pm\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2}\right],$
@@
has correct asymptotic coverage, i.e. it satisfies
$\liminf_{\min(n,N)\to\infty}\Pr(\theta^{\star}\in\mathcal{C}_{\alpha}^{\mathrm{\texttt{FABPP}}})=\liminf_{\min(n,N)\to\infty}\Pr(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f}))\geq 1-\alpha.$
@@
By Section S3.1 (Gaussian prior) and Section S3.1 (horseshoe prior), $\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})$ is an asymptotically valid $1-\delta$ confidence region for $\Delta_{\theta^{\star}}$, that is
$\liminf_{\min(n,N)\to\infty}\text{Pr}(\Delta_{\theta^{\star}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}))$ $\geq 1-\delta.$
@@
$\liminf_{\min(n,N)\to\infty}\text{Pr}(m_{\theta^{\star}}\in\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f}))$ $\geq 1-(\alpha-\delta).$
@@
$E=\{\Delta_{\theta^{\star}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})\}\cap\{m_{\theta^{\star}}\in\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f})\}.$
@@
$\liminf_{n,N\to\infty}\mathrm{Pr}(E)$ $\geq 1-\limsup_{\min(n,N)\to\infty}\mathrm{Pr}(\{\Delta_{\theta^{\star}}\notin\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})\}\cup\{m_{\theta^{\star}}\notin\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f})\})$
$\geq 1-\limsup_{\min(n,N)}\mathrm{Pr}(\{\Delta_{\theta^{\star}}\notin\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})\})-\limsup_{N\to\infty}\mathrm{Pr}(\{m_{\theta^{\star}}\notin\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f})\})$
@@
$0=\Delta_{\theta^{\star}}+m_{\theta^{\star}}\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f}),$
@@
$\liminf_{\min(n,N)\to\infty}\text{Pr}(0\in\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\widehat{\Delta}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}})+\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta^{\star}};\widehat{\sigma}_{\theta^{\star}}^{f}))\geq 1-\alpha,$
@@
S3.2 Section 4.4 - Robustness of FAB-PPI under the Horseshoe Prior
@@
We first state a corollary of Cortinovis & Caron (2024, Theorem 3.4), which follows from the power-law tails of the marginal likelihood under the horseshoe prior (see Section S1.1). The corollary states that, if $|\widehat{\Delta}_{\theta}|$ is very large, then the standard CLT-based confidence interval, $\widehat{\Delta}_{\theta}\pm\widehat{\sigma}_{\theta}z_{1-\delta/2}]$, is recovered.
@@
$\lim_{\Delta\to\pm\infty}\mathcal{R}^{\mathrm{\texttt{FABPP}}}_{\delta}(\Delta;\sigma)-\Delta=[-\sigma z_{1-\delta/2},\sigma z_{1-\delta/2}],$
where the convergence is with respect to the Hausdorff distance on closed subsets of $\mathbb{R}$.
@@
$\mathcal{T}_{\alpha-\delta}(\widehat{m}_{\theta};\widehat{\sigma}_{\theta}^{f})=\left[\widehat{m}_{\theta}-\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2},\widehat{m}_{\theta}+\widehat{\sigma}_{\theta}^{f}z_{1-(\alpha-\delta)/2}\right]$
@@
$\lim_{\Delta\to\pm\infty}\mathcal{S}_{\alpha,\delta}(\Delta,\sigma,m,\sigma^{f})-(\Delta+m)$ $=[-\sigma z_{1-\delta/2}-\sigma^{f}z_{1-(\alpha-\delta)/2},\sigma z_{1-\delta/2}+\sigma^{f}z_{1-(\alpha-\delta)/2}],$
where, again, the convergence is with respect to the Hausdorff distance on closed subsets of $\mathbb{R}$. Therefore, if $|\widehat{\Delta}_{\theta}|\gg 0$, the confidence region $\mathcal{S}_{\alpha,\delta}(\widehat{\Delta}_{\theta},\widehat{\sigma}_{\theta},\widehat{m}_{\theta},\widehat{\sigma}_{\theta}^{f})$ reverts to the standard interval
$\widehat{\Delta}_{\theta}+\widehat{m}_{\theta}\pm(\widehat{\sigma}_{\theta}z_{1-\delta/2}+\widehat{\sigma}^{f}_{\theta}z_{1-(\alpha-\delta)/2})].$
It follows that, if $\inf_{\theta^{\prime}\in\mathbb{R}}|\widehat{\Delta}_{\theta^{\prime}}|\gg 0$, the confidence region for $\theta^{\star}$, defined as
@@
S3.3 Section 4.5 - Consistency of FAB-PPI Mean Estimators
@@
where we recall that $\ell(y;\sigma,\tau)=\log\int_{\mathbb{R}}\mathcal{N}(y;\Delta,\sigma^{2})\pi_{0}(\Delta;\tau)d\Delta$, where $\tau$ is a scale parameter of the prior $\pi_{0}$. By assumption, both PPI estimators, $\widehat{\theta}^{\mathrm{\texttt{PP}}}$ and $\widehat{\theta}^{\mathrm{\texttt{PP+}}}$, are strongly consistent estimators of $\theta^{\star}$. It remains to prove that
@@
Under the horseshoe prior (23), $\ell_{\text{HS}}^{\prime}(y;1,1)$ is bounded. Therefore, (S41) and (S42) hold almost surely by sandwiching.
@@
as $\min(n,N)\to\infty$, where $\widehat{\Sigma}_{\theta}$ is an estimator of $\mathrm{cov}(\widehat{\Delta}_{\theta})$. Again, this holds for both the PPI and PPI++ estimators (Angelopoulos et al., 2023a, b). We consider $d$ independent priors, $\pi_{0}(\Delta_{\theta,k};\widehat{\sigma}_{\theta,k})$ for $k=1,\dots,d$, on the components of $\Delta_{\theta}$, where $\widehat{\sigma}_{\theta,k}^{2}$ is the $k$-th diagonal element of $\widehat{\Sigma}_{\theta}$. The multivariate FAB-PPI estimator $\widehat{\Delta}_{\theta}^{\mathrm{\texttt{FABPP}}}$ is formed by stacking the individual estimators
$\widehat{\Delta}_{\theta,k}^{\mathrm{\texttt{FABPP}}}=\widehat{\Delta}_{\theta,k}+\widehat{\sigma}_{\theta,k}^{2}\ell^{\prime}\left(\widehat{\Delta}_{\theta,k};\widehat{\sigma}_{\theta,k},\widehat{\sigma}_{\theta,k}\right)$
@@
exactly as in the univariate case. Moreover, also multivariate FAB-PPI enjoys asymptotic coverage as $\min(n,N)\to\infty$. In particular, Theorem 4.1 can be easily extended to the multivariate case by applying a union bound over the dimensions of $\Delta_{\theta}$.
@@
Here we provide a brief description of each dataset used for the real data experiments in Section 5.2. For additional details, the reader may refer to Angelopoulos et al. (2023a). All of the datasets were downloaded from the examples provided as part of the ppi-py package (Angelopoulos et al., 2023b).